# From Fibers to Cells: Fourier-Based Registration Enables Virtual Cresyl Violet Staining From 3D Polarized Light Imaging

### v1

@@ 3.1.1 Test data
We manually select four ROIs for model evaluation, ensuring a diverse representation of different cytoarchitectonic characteristics from the held-out test section (Fig. 4B). The ROIs contain primary motor cortex 4a, temporal cortical area TE, the hippocampal cornu Ammonis (CA) region, and parts of the putamen and globus pallidus (Pars interna and Pars externa) as subcortical structures (Fig. 5).
As the computation of image metrics requires a precise alignment of test data, we perform elastic registration of test ROIs based on landmarks and image intensity using the bUnwarpJ Arganda-Carreras et al., (2006) algorithm. We use predictions of an independently trained Gram+Reg model as target. The predictions are used to manually identify 15-25 characteristic cell clusters as landmarks per ROI, which are confirmed by the location and arrangement of faint shadows of cells visible in the 3D-PLI transmittance. For registration, we use an image weight of 1.0, a landmark weight of 10.0, and a consistency weight of 10.0. The strong weights for landmarks and consistency are chosen to prevent the transformation field from overly conforming to the predictions and to overcome local optima.
Figure 5: Overview of ROIs used for the evaluation, which represent distinct cellular architectures. They were extracted from section 559, located 0.6 mm apart from the training sections. Embedded windows show magnified details inside each ROI. The locations of the magnified views are indicated by yellow frames. Columns each show one of the four test ROIs taken from the anterior subdivision of the primary motor cortex (4a), the hippocampal cornu Ammonis (CA) region, temporal cortical area TE, and parts of the putamen and globus pallidus (GP; a: GP Pars interna; b: GP Pars externa) as subcortical nuclei. The first three rows demonstrate 3D-PLI modalities transmittance $I_{T}$, retardation $\sin\delta$ (scaled using gamma correction for visualization) and fiber orientation in HSV color space (hue: direction $\varphi$; saturation, brightness: retardation $\sin\delta$). The 3D-PLI modalities are compared to the registered target Cresyl violet and predicted virtual staining.
@@ 3.1.2 Evaluation scores
We evaluate the impact of different model parameter choices on the quality of the predicted virtual staining by applying structural similarity index measure (SSIM, Wang et al., 2004), mutual information (MI), and root-mean-square error (RMSE). For each metric, we report the mean over all test ROIs. To evaluate the ability to localize cell bodies, we compute F1 scores for instance segmentations of predicted and target cells using a contour proposal network (CPN, Upschulte et al., 2022, 2023). We only consider cells with a minimum in-plane size of 100 µm² and require a minimum intersection over union (IoU) of 30%. Smaller cells could not be reconstructed accurately and would introduce noise into the evaluation. For the cell detection model, we fine-tuned a pre-trained222ginoro_CpnResNeXt101UNet-fbe875f1a3e5ce2c CPN for cell body segmentation in cell-stained microscopy images using the celldetection [^3] https://github.com/FZJ-INM1-BDA/celldetection Python package. Fine-tuning was performed on a diverse mix of manually annotated images, as well as synthetic data.
@@ 3.2 Online registration metric
Figure 6: Robustness of different online registration metrics against synthetic image distortions. We report the proportion of correct rigid alignments for different blur (A) and noise (B) levels by registration metrics MSE, CC, PC, and BIPC. The registration is considered successful (a hit) if the translational displacement obtained remains within 5 pixels of the actual displacement.
To compute a pixel-aligned reconstruction loss we apply the online registration head (Sec. 2.3). We consider correlation-based registration metrics CC and MSE with phase correlation (PC, Kuglin, 1975) and blur-invariant phase correlation (BIPC, Ojansivu and Heikkila, 2007). We apply a Hann window to PC and BIPC before their calculation to mitigate their bias towards the sharp image edges Gonzales and Woods, (2008).
@@
Fig. 6 shows that as blur increases, BIPC with a Hann window achieves the best performance. This result is expected, as the metric remains invariant to blur. We also observe that CC and MSE perform best in mitigating the effects of noise, while the other metrics fail already with a minor amount of noise. Overall, MSE provides a balanced tradeoff between blur and noise response.
Table 1: Performance of models trained with different online registration metrics in terms of similarity of generated images with ground truth. We compare cross-correlation (CC), phase-correlation (PC), blur-invariant phase-correlation(BIPC), and mean-squared error (MSE) as online registration metrics against using no online registration (-).
@@
Gram - 0.126 33.6 0.354 48.8
Gram+Reg CC 0.136 33.0 0.379 55.2
BIPC 0.149 31.5 0.371 54.0
PC 0.185 31.9 0.415 59.4
MSE 0.226 29.8 0.444 63.7
@@
Tab. 1 shows a quantitative comparison of the models. Independent of the metric, the best results are observed when using Gram+Reg with online registration Among all metrics considered, the model with MSE performs best.
@@ 3.3 Performance analysis of different model variants
Figure 7: Including equivariance loss $\mathcal{L}_{E}$ in the training improves cell instance overlap of Gram+Reg and GAN+Reg models. (A) 3D-PLI input as transmittance $I_{T}$, retardation $\sin\delta$ (scaled using gamma correction) and fiber orientation map in HSV color space (hue: direction $\varphi$; saturation, brightness: retardation $\sin\delta$) for an example patch from the pyramidal layer of the hippocampal CA1 region. (B) Registered target Cresyl violet staining, a cell segmentation by a CPN model, and the identified target cell instances. (C) Predicted virtual staining of models trained with equivariance loss $\mathcal{L}_{E}$ enabled (✓) or disabled (✗). Cell instances are overlaid with their target cells. Models trained with equivariance loss show less displacement of cells and increased F1 scores for this specific image patch. Since the displacement occurs randomly, a patch and random seeds for the models are selected that show a particularly strong effect.
To justify including equivariance loss $\mathcal{L}_{E}$ (Eq. 9) in the total loss, Fig. 7 illustrates the effect of training Gram+Reg models with this loss component enabled and disabled. With $\mathcal{L}_{E}$ disabled ($\eta$ = 0), cell instances in the model predictions exhibit a relative displacement. Enabling $\mathcal{L}_{E}$ ($\eta$ = 0.1) improves the overlap of cell instances with the original Cresyl violet staining.
Table 2: Effect of the proposed online registration on models trained with different choices of style loss $\mathcal{L}_{S}$. Each variant is trained with online registration head (Reg.) and equivariance loss ($\mathcal{L}_{E}$) enabled or disabled. MSE metric is used for online registration and $\eta$ = 0.1 for weighting of $\mathcal{L}_{E}$. The deviation is reported as standard error over four independent trainings with different random seeds.
@@
To quantify the impact of the loss component $\mathcal{L}_{E}$ on the predictions, we compare the models with $\mathcal{L}_{E}$ enabled and disabled in Tab. 2. We consider Gram and Gram+Reg, as well as GAN and GAN+Reg as model variants. All models use $\lambda=0.5$ expect for the Gram model without online registration, where $\lambda=0.1$ showed better performance. Since displacements of cell instances in the model predictions occur randomly with varying strength, we report mean and standard errors across four independent trainings. Models trained with $\mathcal{L}_{E}$ enabled show the highest correspondence with the target staining. Without online registration, the advantage of including $\mathcal{L}_{E}$ becomes negligible. Furthermore, results in Tab. 2 demonstrate that Gram models outperform their corresponding GAN models in all evaluation metrics. Regardless of the variant used for style loss $\mathcal{L}_{S}$, all models benefit from the online registration.
@@
Figure 8: Comparison of virtual stainings for a patch of temporal area TE. (A) 3D-PLI input visualized as fiber orientation in HSV color space (hue: direction $\varphi$; saturation, brightness: retardation $\sin\delta$) and transmittance $I_{T}$. (B) Registered target Cresyl violet staining for reference with annotations of cortical layers I-VI and white matter (wm). (C) Predicted virtual stainings of GAN and Gram models and extended variants using MSE online registration (GAN+Reg, Gram+Reg). (D) GLI images ranging from the pial surface to the cortex/white matter transition and average profiles (red lines) for layers II - VI.
@@
A comparison of model predictions is shown in Fig. 8C for the whole cortical depth of an Isocortex sample from temporal cortical area TE. GAN+Reg and Gram+Reg provide the most realistic-looking reconstructions in terms of relative size and shape of cell bodies, along with the differences between layers concerning cell packing density. While Gram+Reg seems to be influenced by technically-related inhomogeneities in staining intensity, GAN+Reg produces a clearer contrast between stained cells and surrounding tissue. Gram+Reg detects the transition from the cortex to white matter better than GAN+Reg. Both methods introduce an artificial arrangement of cells into cortical columns not present in the original Cresyl violet staining.
@@
To further analyze the ability of the methods to reconstruct the laminar cell organization, we compute grey level index (GLI) values Schleicher et al., (1978); Zilles et al., (1978). GLI values provide an established proxy for volume density of stained cell bodies in gray matter regions and are used to characterize the laminar architecture of cortical areas Schleicher et al., (2000). To compute GLI values, adaptive thresholding is applied to create a mask of pixels occupied by cell bodies for each image. The masks are subsequently down scaled by a factor of 16 to a spatial resolution of 20.8 µm, with each value representing the fraction of segmented pixels. Resulting GLI images, cropped to the area between pial surface and the cortex/white matter transition, are displayed for the Cresyl violet target and the predictions in Fig. 8D.
@@
For each of the GLI images, 31 intensity profiles vertically oriented to the cortical layers are extracted to represent the columnar distribution of GLI values. The profiles are restricted to layers II - VI since staining inhomogeneities make GLI values for layer I unreliable. To obtain one representative profile for each image, the profiles are averaged and smoothed using a mean filter with kernel size 3 to improve the signal-to-noise ratio. The average profiles are shown on top of each GLI image in Fig. 8D.
@@
3.4 Optimal balancing of style and content
Table 3: Different values for weighting parameter $\lambda$ show that a balance between style loss $\mathcal{L}_{S}$ and reconstruction loss $\mathcal{L}_{R}$ is required in the training of a Gram+Reg model. Setting $\lambda=0$ means only $\mathcal{L}_{S}$ is used. Setting $\lambda=1$ means only $\mathcal{L}_{R}$ is used.
@@ 3.4 Optimal balancing of style and content
Figure 9: Predicted virtual Cresyl violet stainings for different weightings between reconstruction and style loss components. (A) 3D-PLI inputs as transmittance $I_{T}$, retardation $\sin\delta$ and fiber orientation in HSV color space (hue: direction $\varphi$; saturation, brightness: retardation $\sin\delta$). The retardation has been scaled using gamma correction for visualization. (B) Registered target Cresyl violet staining. (C) Predicted virtual stainings of models trained with different weightings $\lambda$ between reconstruction loss $\mathcal{L}_{R}$ and style loss $\mathcal{L}_{S}$. Setting $\lambda=0$ focuses on $\mathcal{L}_{S}$ and setting $\lambda=1$ on $\mathcal{L}_{R}$ exclusively. Gram loss is used for $\mathcal{L}_{S}$. Within each row, yellow markers indicate the same structures across all columns. (i) Temporal area TE and underlying white matter with arrowheads highlighting blood vessels within white matter and cortical layer VI. (ii) Dentate gyrus of the hippocampus with a dashed line delineating the proximal end of the granular layer of the fascia dentata (FD). (iii) Pyramidal layer of the cornu Ammonis (CA1) region of the hippocampus highlighted by dashed lines. (iv) Layer V of the primary motor cortex with arrowheads highlighting two Betz cells.
@@
To identify an optimal weighting $\lambda$ of the style loss in Eq. 10, we train Gram+Reg models with different values for $\lambda$ and MSE online registration metric. A quantitative evaluation of the trained models in Tab. 3 shows that larger values for $\lambda$ up to $\lambda$ = 0.75 achieve the best MI, RMSE, and SSIM scores. This indicates that the reconstruction loss should be given a stronger weight but should not be used alone ($\lambda=1$). The highest overlap of predicted cell instances measured by the F1 score is obtained for a balanced weighting with $\lambda$ = 0.5.
@@
A visual comparison of predictions of Gram+Reg models trained with different choices for $\lambda$ is shown in Fig. 9 for selected crops from the test data. By using $\mathcal{L_{R}}$ only (i.e. $\lambda=1$), the model is unable to reconstruct details from the Cresyl violet staining. It is only able to locate strongly pronounced structures such as the granular layer delineated in (ii). Independent of the style loss weighting, the models do not resolve individual cell instances present within the granular layer. Instead, they capture the overall cell density, represented by a continuous dark purple color. By increasing emphasis on $\mathcal{L}_{S}$ (i.e. with decreasing values for $\lambda$), the generator is able to reconstruct details such as neuronal cell bodies in less dense regions in (i)-(iv) and blood vessels in (i). It is also able to generate Glial cells to match the appearance of white matter in (i) and (iii). However, generators trained with an overweighting of style loss also tend to miss some strongly pronounced structures in the predictions, such as the blood vessels shown in (i) or the Betz cells in (iv).
3.5 Reliability of predicted cell instances across cell sizes

### v2

@@ 3.1.1 Test data
We manually select four ROIs for model evaluation, ensuring a diverse representation of different cytoarchitectonic characteristics from the held-out test section (Fig.˜4B). The ROIs contain primary motor cortex 4a, temporal cortical area TE, the hippocampal cornu Ammonis (CA) region, and parts of the putamen and globus pallidus (Pars interna and Pars externa) as subcortical structures (Fig.˜5).
As the computation of image metrics requires a precise alignment of test data, we perform elastic registration of test ROIs based on landmarks and image intensity using the bUnwarpJ arganda-carreras2006 algorithm. We use predictions of an independently trained Gram+Reg model as target. The predictions are used to manually identify 15-25 characteristic cell clusters as landmarks per ROI, which are confirmed by the location and arrangement of faint shadows of cells visible in the 3D-PLI transmittance. For registration, we use an image weight of 1.0, a landmark weight of 10.0, and a consistency weight of 10.0. The strong weights for landmarks and consistency are chosen to prevent the transformation field from overly conforming to the predictions and to overcome local optima.
Figure 5: Overview of ROIs used for the evaluation, which represent distinct cellular architectures. They were extracted from section 559, located 0.6 mm apart from the training sections. Embedded windows show magnified details inside each ROI. Columns each show one of the four test ROIs taken from the anterior subdivision of the primary motor cortex (4a), the hippocampal cornu Ammonis (CA) region, temporal cortical area TE, and parts of the putamen and globus pallidus (GP; a: GP Pars interna; b: GP Pars externa) as subcortical nuclei. The first three rows demonstrate 3D-PLI modalities transmittance, retardation (scaled using gamma correction for visualization) and fiber orientation in HSV color space (hue: fiber direction; saturation/brightness: retardation). The 3D-PLI modalities are compared to the registered target Cresyl violet and predicted virtual staining.
@@ 3.1.2 Evaluation scores
We evaluate the impact of different model parameter choices on the quality of the predicted virtual staining by applying structural similarity index measure (SSIM, wang2004), mutual information (MI), and root-mean-square error (RMSE). For each metric, we report the mean over all test ROIs. To evaluate how well cell positions are preserved by different models, we compute F1 scores based on cell instance segmentations by a contour proposal network (CPN, upschulte2022; upschulte2023). Predicted cell instances are obtained from segmentations of the virtual staining and compared to target cell instances from the corresponding Cresyl violet images. For each ROI, predicted and target cells are matched by calculating their intersection over union (IoU), requiring a minimum IoU of 30% for a match. Each target cell can be a match for at most one predicted cell. Matched instances are counted as true positives, unmatched predicted cells as false positives, and unmatched target cells as false negatives. F1 scores are then computed from aggregated counts across all ROIs. For the cell detection model, we fine-tuned a pre-trained CPN222ginoro_CpnResNeXt101UNet-fbe875f1a3e5ce2c for cell body segmentation in cell-stained microscopy images using the celldetection [^3] https://github.com/FZJ-INM1-BDA/celldetection Python package. Fine-tuning was performed on a diverse mix of manually annotated images, as well as synthetic data.
@@ 3.2 Online registration metric
Figure 6: Robustness of different online registration metrics against synthetic image distortions. We report the proportion of correct rigid alignments for different blur (A) and noise (B) levels by registration metrics MSE, $C$, PC, and BIPC. The registration is considered successful (a hit) if the translational displacement obtained remains within 5 pixels of the actual displacement.
To compute a pixel-aligned reconstruction loss we apply the online registration head (Sec.˜2.3). We consider correlation-based registration metrics CC and MSE with phase correlation (PC, kuglin1975) and blur-invariant phase correlation (BIPC, ojansivu2007). We apply a Hann window to PC and BIPC before their calculation to mitigate their bias towards the sharp image edges gonzalez2008digital.
@@
Fig.˜6 shows that as blur increases, BIPC with a Hann window achieves the best performance. This result is expected, as the metric remains invariant to blur. We also observe that CC and MSE perform best in mitigating the effects of noise, while the other metrics fail already with a minor amount of noise. Overall, MSE provides a balanced tradeoff between blur and noise response.
Table 1: Performance of models trained with different online registration metrics in terms of similarity of generated images with ground truth. We compare cross-correlation ($C$), phase-correlation (PC), blur-invariant phase-correlation(BIPC), and mean-squared error (MSE) as online registration metrics against using no online registration (-). Arrows indicate the direction of better performance ($\uparrow$ higher is better, $\downarrow$ lower is better). Best scores per column in bold.
@@
Gram - 0.126 33.6 0.354 30.2
Gram+Reg $C$ 0.136 33.0 0.379 34.4
BIPC 0.149 31.5 0.371 32.6
PC 0.185 31.9 0.415 38.0
MSE 0.226 29.8 0.444 41.3
@@
Tab.˜1 shows a quantitative comparison of the models. Independent of the metric, the best results are observed when using Gram+Reg with online registration Among all metrics considered, the model with MSE performs best.
@@ 3.3 Performance analysis of different model variants
Figure 7: Example image illustrating how equivariance loss $\mathcal{L}_{E}$ improves cell instance overlap of Gram+Reg and GAN+Reg models. (A) 3D-PLI input as transmittance, retardation (scaled using gamma correction) and fiber orientation map in HSV color space (hue: fiber direction; saturation/brightness: retardation). (B) Registered target Cresyl violet staining, a cell segmentation by a CPN model, and the identified target cell instances. (C) Predicted virtual staining of models trained with equivariance loss $\mathcal{L}_{E}$ enabled (✓) or disabled (✗). Predicted cell instances (red) by each model are overlaid with the target cells (gray) and F1 scores computed for cells in this specific image patch. Since the displacement occurs randomly, a patch from our test set and random seeds for the models were manually selected that illustrate that effect. The patch shows the pyramidal layer of the hippocampal CA1 region.
To justify including equivariance loss $\mathcal{L}_{E}$ (Eq.˜9) in the total loss, Fig.˜7 illustrates the effect of training different models with this loss component enabled and disabled. With $\mathcal{L}_{E}$ disabled ($\eta$ = 0), cell instances in the model predictions exhibit a relative displacement. Enabling $\mathcal{L}_{E}$ ($\eta$ = 0.1) improves the overlap of cell instances with the original Cresyl violet staining.
Table 2: Effect of the proposed online registration on models trained with different choices of style loss $\mathcal{L}_{S}$. Each variant is trained with online registration head (Reg.) and equivariance loss ($\mathcal{L}_{E}$) enabled or disabled. MSE metric is used for online registration and $\eta$ = 0.1 for weighting of $\mathcal{L}_{E}$. The deviation is reported as standard error over four independent trainings with different random seeds. Arrows indicate the direction of better performance ($\uparrow$ higher is better, $\downarrow$ lower is better). Best scores per column in bold.
@@
To quantify the impact of the loss component $\mathcal{L}_{E}$ on the predictions, we compare the models with $\mathcal{L}_{E}$ enabled and disabled in Tab.˜2. We consider Gram and Gram+Reg, as well as GAN and GAN+Reg as model variants. All models use $\lambda=0.5$ expect for the Gram model without online registration, where $\lambda=0.1$ showed better performance. Since displacements of cell instances in the model predictions occur randomly with varying strength, we report mean and standard errors across four independent trainings. Models trained with $\mathcal{L}_{E}$ enabled show the highest correspondence with the target staining. Without online registration, the advantage of including $\mathcal{L}_{E}$ becomes negligible. Furthermore, results in Tab.˜2 demonstrate that Gram models outperform their corresponding GAN models in all evaluation metrics. Regardless of the variant used for style loss $\mathcal{L}_{S}$, all models benefit from the online registration. We present averaged scores across all test ROIs. Individual scores per ROI, are provided in Tab.˜5 in the appendix as additional information.
@@
Figure 8: Comparison of virtual stainings for a patch of temporal area TE. (A) 3D-PLI input visualized as fiber orientation in HSV color space (hue: fiber direction; saturation/brightness: retardation) and transmittance. (B) Registered target Cresyl violet staining for reference with annotations of cortical layers I-VI and white matter (wm). (C) Predicted virtual stainings of GAN and Gram models and extended variants using MSE online registration (GAN+Reg, Gram+Reg). (D) GLI images ranging from the pial surface to the cortex/white matter transition and average profiles (red lines) for layers II - VI.
@@
A comparison of model predictions is shown in Fig.˜8C for the whole cortical depth of an Isocortex sample from temporal cortical area TE. GAN+Reg and Gram+Reg provide the most realistic-looking reconstructions in terms of relative size and shape of cell bodies, along with the differences between layers concerning cell packing density. While Gram+Reg seems to be influenced by technically-related inhomogeneities in staining intensity, GAN+Reg produces a clearer contrast between stained cells and surrounding tissue. Gram+Reg detects the transition from the cortex to white matter better than GAN+Reg. Both methods introduce an artificial arrangement of cells into cortical columns not present in the original Cresyl violet staining.
@@
To further analyze the ability of the methods to reconstruct the laminar cell organization, we compute grey level index (GLI) values zilles1978. GLI values provide an established proxy for volume density of stained cell bodies in gray matter regions and are used to characterize the laminar architecture of cortical areas schleicher2000. To compute GLI values, adaptive thresholding is applied to create a mask of pixels occupied by cell bodies for each image. The masks are subsequently down scaled by a factor of 16 to a spatial resolution of 20.8 µm, with each value representing the fraction of segmented pixels. Resulting GLI images, cropped to the area between pial surface and the cortex/white matter transition, are displayed for the Cresyl violet target and the predictions in Fig.˜8D.
@@
For each of the GLI images, 31 intensity profiles vertically oriented to the cortical layers are extracted to represent the columnar distribution of GLI values. The profiles are restricted to layers II - VI since staining inhomogeneities make GLI values for layer I unreliable. To obtain one representative profile for each image, the profiles are averaged and smoothed using a mean filter with kernel size 3 to improve the signal-to-noise ratio. The average profiles are shown on top of each GLI image in Fig.˜8D.
@@
3.4 Effect of loss weighting parameters
Table 3: Ablation study on weighting parameter $\eta$ of equivariance loss $\mathcal{L}_{E}$ for a Gram+Reg model. Performance is robust across a wide range of $\eta>0$ values. Only when the loss term is completely removed ($\eta=0$), we observe a measurable degradation in performance. The deviation is reported as standard error over four independent trainings with different random seeds. Arrows indicate the direction of better performance ($\uparrow$ higher is better, $\downarrow$ lower is better). Best scores per column in bold.
Method $\eta$ MI $\uparrow$ RMSE $\downarrow$ SSIM $\uparrow$ F1 $\uparrow$
Gram+Reg 0.0 0.211 $\pm$ 0.004 30.2 $\pm$ 0.4 0.432 $\pm$ 0.004 35.2 $\pm$ 1.2
0.01 0.225 $\pm$ 0.002 29.5 $\pm$ 0.3 0.446 $\pm$ 0.001 41.6 $\pm$ 0.2
0.1 0.224 $\pm$ 0.003 29.6 $\pm$ 0.6 0.445 $\pm$ 0.003 41.0 $\pm$ 0.4
1.0 0.220 $\pm$ 0.001 29.4 $\pm$ 0.1 0.445 $\pm$ 0.001 41.4 $\pm$ 0.2
10.0 0.224 $\pm$ 0.002 29.4 $\pm$ 0.3 0.446 $\pm$ 0.001 41.2 $\pm$ 0.2
To incorporate equivariance loss $\mathcal{L}_{E}$ into previous models, we set its weighting parameter to $\eta=0.1$. To validate this choice, we perform training of Gram+Reg models with $\eta$ values ranging from 0 to 10. Results in Tab.˜3 indicate that performance is robust for $\eta>0$. Only when the equivariance loss is entirely removed from training (setting $\eta=0$), we observe a measurable negative effect on the results. We report mean values and standard errors over four independent trainings to show the significance of this effect.
Table 4: Different values for weighting parameter $\lambda$ show that a balance between style loss $\mathcal{L}_{S}$ and reconstruction loss $\mathcal{L}_{R}$ is required in the training of a Gram+Reg model. Setting $\lambda=0$ means only $\mathcal{L}_{S}$ is used. Setting $\lambda=1$ means only $\mathcal{L}_{R}$ is used. Arrows indicate the direction of better performance ($\uparrow$ higher is better, $\downarrow$ lower is better). Best scores per column in bold.
@@ 3.4 Effect of loss weighting parameters
Figure 9: Predicted virtual Cresyl violet stainings for different weightings between reconstruction and style loss components. (A) 3D-PLI inputs as transmittance, retardation and fiber orientation in HSV color space (hue: fiber direction; saturation/brightness: retardation). The retardation has been scaled using gamma correction for visualization. (B) Registered target Cresyl violet staining. (C) Predicted virtual stainings of models trained with different weightings $\lambda$ between reconstruction loss $\mathcal{L}_{R}$ and style loss $\mathcal{L}_{S}$. Setting $\lambda=0$ focuses on $\mathcal{L}_{S}$ and setting $\lambda=1$ on $\mathcal{L}_{R}$ exclusively. Gram loss is used for $\mathcal{L}_{S}$. Within each row, yellow markers indicate the same structures across all columns. (i) Temporal area TE and underlying white matter with arrowheads highlighting blood vessels within white matter and cortical layer VI. (ii) Dentate gyrus of the hippocampus with a dashed line delineating the proximal end of the granular layer of the fascia dentata (FD). (iii) Pyramidal layer of the cornu Ammonis (CA1) region of the hippocampus highlighted by dashed lines. (iv) Layer V of the primary motor cortex with arrowheads highlighting two Betz cells.
@@
To identify an optimal weighting $\lambda$ of the style loss in Eq.˜10, we train Gram+Reg models with different values for $\lambda$ and MSE online registration metric. A quantitative evaluation of the trained models in Tab.˜4 shows that larger values for $\lambda$ up to $\lambda$ = 0.75 achieve the best evaluation scores. This indicates that the reconstruction loss should be given a stronger weight but should not be used exclusively ($\lambda=1$).
@@
A visual comparison of predictions of Gram+Reg models trained with different choices for $\lambda$ is shown in Fig.˜9 for selected crops from the test data. By using $\mathcal{L_{R}}$ only (i.e. $\lambda=1$), the model is unable to reconstruct details from the Cresyl violet staining. It is only able to locate strongly pronounced structures such as the granular layer delineated in (ii). Independent of the style loss weighting, the models do not resolve individual cell instances present within the granular layer. Instead, they capture the overall cell density, represented by a continuous dark purple color. By increasing emphasis on $\mathcal{L}_{S}$ (i.e. with decreasing values for $\lambda$), the generator is able to reconstruct details such as neuronal cell bodies in less dense regions in (i)-(iv) and blood vessels in (i). It is also able to generate Glial cells to match the appearance of white matter in (i) and (iii). However, generators trained with an overweighting of style loss also tend to miss some strongly pronounced structures in the predictions, such as the blood vessels shown in (i) or the Betz cells in (iv).
3.5 Influencing factors on model predictions
To assess the reliability of our method under varying biological and imaging conditions, we analyze in how far the reconstruction of cell instances depends on cell size and local strength of birefringence. We also examine how different focus levels of the LMP-1 impair the quality of the virtual staining.